Noel has $350,000 with which to purchase an ordinary annuity delivering monthly payments for 20 years after a 10-year period of deferral. What monthly payment will he receive, if the undistributed funds earn 5.5% compounded semiannually?

  Monthly payment $
Answer #1

Aus Possent valul (PV) $350,000 g iven rate i 5.51. of comprended Seniannually su5 -- 2 2:75.1. 0.0275 Tafal in zo glas NumbeWe know that, man PU=CC-11+8) (1-(1+rjMj g g - (men) Pv= (1-(*+r) I+ (1+07] Pv=ic (@to jilm th] 11+8) + after Subhi faling, 3C %3D 2 49957.5०५ hence, we get monthly Payment - in $249952. 5०५